OpenStudy (anonymous):

Hi, can you look at this problem? squareroot X+5 - squareroot x-3 = 4

OpenStudy (ribhu):

yeah \[\sqrt{x+5}-\sqrt{x-3}= 4\]

OpenStudy (anonymous):

yes, How should I start off?

OpenStudy (ribhu):

take sqrt x-3 to rhs, then square both the sides

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

\[x+5= 4/x+9\]

OpenStudy (ribhu):

x+5 = (4 + sqrt x-3)^2 not that u wrote.

OpenStudy (anonymous):

ops lol

OpenStudy (anonymous):

x + 5 = 7 +x

OpenStudy (anonymous):

2x = 12

OpenStudy (anonymous):

x = 6?

OpenStudy (ribhu):

no

OpenStudy (anonymous):

ops

OpenStudy (anonymous):

x+5 = 16+3+x

OpenStudy (ribhu):

x+5 = 16+x-3+8sqrt x-3

OpenStudy (ribhu):

no what r u doin

OpenStudy (anonymous):

x+5= 4^2 + squareroot x -3

OpenStudy (anonymous):

x+5 = 16 +x+3

OpenStudy (ribhu):

no its that what i wrote.

OpenStudy (ribhu):

x= 4

OpenStudy (anonymous):

but squarerrot (x-3)^2 = x+3

OpenStudy (anonymous):

how do you get what you got?

OpenStudy (ribhu):

no its x-3

OpenStudy (anonymous):

my bad

OpenStudy (ribhu):

yeah.

OpenStudy (anonymous):

ok so x+5=16+x-3

OpenStudy (ribhu):

and+8sqrtx-3

OpenStudy (ribhu):

on the rhs

OpenStudy (anonymous):

how? do you get that?